### 22 Dec 2023 v5

@@ 1 Introduction
Astronomy often involves studying objects so distant that they remain unresolved point sources with even the largest telescopes. Exoplanets are particularly difficult to study because they are small compared to other astronomical objects. Short of building a large interferometer to spatially resolve the disks of exoplanets, we must rely on chance and astronomical trickery to map their atmospheres and surfaces. [^1] The diffraction limit dictates that resolving the disk of an Earth-sized planet at 10 pc requires a telescope—or array of telescopes—at least 25 km across at optical wavelengths, or 500 km across in the thermal infrared.
@@ 2 Mapping Basics
We focus on the analysis of planetary light, be it thermal radiation or reflected starlight. Planetary light can be separated from starlight temporally (Charbonneau et al. 2005; Deming et al. 2005), spatially (Marois et al. 2008), or spectrally (Brogi et al. 2012). We also briefly discuss transmission mapping below.
@@ 2.1 Transmission Mapping
It is also possible to glean spatial information about an exoplanet’s atmosphere when it transits in front of its host star: transmission mapping, or transit mapping. We briefly describe transmission mapping, as it shares some concepts with emission and reflection mapping. Transmission mapping relies on two geometrical facts: 1) only part of the planetary limb is backlit during transit ingress and egress, and 2) the planetary limb probed in transmission moves on the planet during a transit.
At the start of transit ingress, only the western (leading) limb of the planet is in front of the star and contributes to the transmission spectrum; likewise, the end of egress is only sensitive to the eastern (trailing) limb of the planet. Such transmission mapping has been proposed for detecting the difference in atmospheric scale height (Dobbs-Dixon et al. 2012), winds (Louden and Wheatley 2015), and aerosols (Kempton et al. 2017; Powell et al. 2019) near the eastern and western terminators. Many codes have been developed to perform such transmission mapping (Espinoza and Jones 2021; MacDonald and Lewis 2022; Grant and Wakeford 2023). The ingress-egress asymmetry is most marked for planets that are large compared to their star; the best spatial resolution could be obtained for a giant planet orbiting a white dwarf (e.g., Vanderburg et al. 2020). This phenomenon is conceptually similar to eclipse mapping, but is very sensitive to the limb-darkening of the star. Transmission mapping has already been performed with ground-based high resolution spectroscopy (HRS; Ehrenreich et al. 2020). Even in the middle of transit, HRS can differentiate between the blue-shifted eastern limb and and the red-shifted western limb (Boucher et al. 2023), a technique analogous to Doppler imaging.
Moreover, the planet will rotate during the transit, i.e., the planetary limb does not exactly coincide with the day–night terminator. In fact, the limb only corresponds to the day–night terminator when the planet is in the center of the stellar disk, which never occurs for a non-zero impact parameter. At the start of transit the limb runs on the dayside of the eastern terminator and on the nightside of the western terminator, and vice versa at the end of transit. The angle between the limb and terminator is proportional to the angular size of the star as seen from the planet, and is therefore most important for ultra short period planets (Nguyen et al. 2020). If a rocky planet has a negligible atmosphere, then topographical features may rotate in and out of view during the transit (McTier and Kipping 2018). Transmission mapping of exoplanets is an exciting growth area, but for the remainder of this review, we focus on mapping exoplanets using light emitted or reflected by the planets themselves.
@@ 3 Mapping Formalism
Arguably the most important simplification we have made is assuming a static map, ${\partial M}/{\partial t}=0$, as it is generally intractable to map the surface of an unresolved planet when that map is changing. [^2] One can map a slowly varying planet, e.g., large-scale cloud patterns on Earth evolve slowly compared to the planet’s rotation (Kawahara and Masuda 2020; Teinturier et al. 2022) and some eccentric planets may still be amenable to eclipse mapping, despite seasonal variations. Beyond these exceptions, one must adopt a parameterized time-variable map, $M(\theta,\phi,t)$. For example, phase curves of eccentric hot Jupiters have been fitted with energy balance models (Lewis et al. 2013; de Wit et al. 2016; Dang et al. 2022). Finally, Cowan et al. (2017) showed that for planets on circular, edge-on orbits one can infer a time-variable map based on the presence of odd harmonics; nonetheless, this is a far cry from mapping the changing surface of a planet. Time-variations in flux therefore come in through the kernel. If there are no occultations, then the changing sub-observer and sub-stellar locations dictate the time-variability of the kernel and hence the time-variations in observed flux.
@@ 3.1 The Inverse Problem
If the orientation of a planet is not known a priori, then the problem is even more challenging because the kernel is a function of one or more unknown parameters: spin frequency, spin orientation, etc. Nonetheless, it has been demonstrated in numerical experiments that one can extract a planet’s spin (rotation rate, obliquity and its direction) from reflected lightcurves (Pallé et al. 2008; Oakley and Cash 2009; Kawahara and Fujii 2010, 2011; Fujii and Kawahara 2012; Schwartz et al. 2016; Kawahara 2016; Farr et al. 2018; Nakagawa et al. 2020) or from thermal lightcurves (Gaidos and Williams 2004; Gómez-Leal et al. 2012; Cowan et al. 2012b, 2013). [^3] A similar formalism has been developed for imaging exoplanets with a solar gravitational lens (Toth and Turyshev 2023).
Exoplanets can be mapped in monochrome using single-band photometry, or using spectrally-resolved data to produce maps at each wavelength. [^4] Mapping exoplanets with ground-based high spectral resolution is another beast entirely: the signal is in the time-variable molecular line shape. We point the interested reader to the excellent review of high resolution atmospheric characterization (Birkby 2018) and note that much work remains to be done in extracting spatial information from high resolution data. It is often more insightful, however, to assume that the lightcurves are correlated due to viewing geometry, molecular absorption, or surface spectra. The inverse problem of mapping a planet using template spectra—let alone retrieving intrinsic surface colors and spectra—based on multi-band photometry or time-resolved spectroscopy is beyond the scope of this review, but has been considered in the literature both in the context of reflected light (Cowan et al. 2009, 2011; Fujii et al. 2010, 2011; Cowan and Strait 2013; Fujii et al. 2017; Kawahara 2020) and thermal emission (Kostov and Apai 2013; Buenzli et al. 2014; Stevenson et al. 2014; Feng et al. 2016; Irwin et al. 2020; Changeat and Al-Refaie 2020; Cubillos et al. 2021).
@@ 4.1 Pixels and Slices
An intuitive approach is to pixelize the planetary surface. The most general pixelization scheme is 2-dimensional, for example, a latitude–longitude grid or Hierarchical Equal Area isoLatitude Pixelization (HEALPix) [^5] http://healpix.sourceforge.net/, but more exotic schemes have been developed for occultation mapping (Louden and Kreidberg 2018). Pixels are useful if the kernel moves or changes its shape in more than one direction, e.g., eclipse mapping or rotational reflected light curves at multiple orbital phases.
@@
For rotational mapping, latitudinal information is difficult to extract, so it is often reasonable to consider a map consisting of uniform slices (like a beach ball). This situation includes rotational mapping in the thermal infrared (Knutson et al. 2007) and rotational mapping of reflected light at a fixed orbital phase (Cowan et al. 2009).
@@ 4.3 The Kernel
Due to the unchanging kernel shape for rotational or orbital thermal mapping, a large fraction of spherical harmonics are in the nullspace, limiting the fidelity of retrieved maps. A common strategy is to ignore basis maps in the nullspace, setting their amplitude to zero in fits (Cowan and Agol 2008). But this strategy can sometimes produce retrieved maps that appear to have regions of negative flux, an artifact of the band-limited deconvolution. The solution is to fit for more basis maps—including those in the nullspace: immeasurable fitted parameters may help the map remain everywhere positive (Keating et al. 2019). [^6] These shenanigans are no help for lightcurves that go negative: that unphysical scenario is a symptom of uncorrected noise (Keating and Cowan 2017; Keating et al. 2019; Beatty et al. 2019). Such hidden variables can be further constrained by a regularization scheme or Gaussian process. Indeed, the most general and robust approach to mapping is likely a high-order orthogonal map expansion—many spherical harmonics or pixels—embedded in a hierarchical model (Farr et al. 2018; Luger et al. 2021). If the prospect of fitting unseen and unknowable parameters by relying on priors is daunting, rest assured that you can often obtain a good fit and reasonable map by adopting an altogether different parameterization, such as slices or pixels (Beatty et al. 2019).
@@
If a planet passes directly behind its star [see chapter on eclipses by Roi Alonso], then one may map its day side. Eclipse mapping can be applied to either planetary emission or reflected light, but the contrast ratio makes the latter daunting. While the eclipse depth is proportional to the integrated day side brightness of the planet, the detailed shape of ingress and egress is a function of the spatial distribution of flux on the planet’s day side (Figure 3). Although eclipses of planets by their host star have so far garnered the most attention, occultations by moons or other planets are also possible, especially in packed planetary systems like TRAPPIST-1 (Veras and Breedt 2017; Luger et al. 2017).
@@ 5 Current Results & Future Prospects
The atmospheres of a few dozen exoplanets have been mapped via their thermal phase curves [^7] The ExoAtmospheres database includes phases curves of exoplanets.. These phase measurements has so far been mostly IRAC photometry from the Spitzer Space Telescope (e.g., Bell et al. 2021; May et al. 2022), or spectroscopy with the Hubble Space Telescope’s WFC3 (e.g., Stevenson et al. 2014; Kreidberg et al. 2018). We send the interested reader to the brief history of thermal phase curves in Charnay et al. (2022). Phase measurements from the James Webb Space Telescope have started being published, but so far focusing primarily on photometry (Mikal-Evans et al. 2023; Bell et al. 2023; Kempton et al. 2023). Rotational phase variations of brown dwarfs have also been measured and converted into brightness maps (e.g., Apai et al. 2017, and chapter by Artigau).
The Ariel mission will measure spectroscopic phase curves for dozens—if not hundreds—of planets and brown dwarfs (Charnay et al. 2022). The wavelength coverage of Ariel and JWST should be able to discriminate between thermal emission and reflected light (Schwartz and Cowan 2015; Parmentier et al. 2016; Keating and Cowan 2017), as well as disentangling phase variations, star spots, Doppler beaming, ellipsoidal variations, and gas accretion (Cowan et al. 2012a; Knutson et al. 2012; Esteves et al. 2015; Dang et al. 2018; Bell et al. 2019).
@@
Ground-based direct imaging could detect variations in thermal emission from directly-imaged planets due to rotation of patchy clouds in and out of view (Biller et al. 2021), as is currently performed for brown dwarfs [see Artigau chapter]. The next generation of ground based telescopes should enable both photometric mapping and Doppler tomography of giant planets (Crossfield et al. 2014), while the Large Interferometer for Exoplanets (Quanz et al. 2022) could map the climates of temperate terrestrial planets (e.g., Gaidos and Williams 2004; Cowan et al. 2012c; Mettler et al. 2023).
@@ Acknowledgements.
  • Robinson, T. : “Detecting Habitability”
  • Alonso, R. : “Characterization of Exoplanets: Secondary Eclipses”
  • Parmentier, V. & Crossfield, I. : “Characterization of Exoplanets: Observations and Modeling of Orbital Phase Curves”
  • Artigau, E. : “Variability of Brown Dwarfs”
  • Birkby, J. : “Spectroscopic Direct Detection of Exoplanets”

### 30 Jan 2024 v6

@@ 1 Introduction
Astronomy often involves studying objects so distant that they remain unresolved point sources with even the largest telescopes. Exoplanets are particularly difficult to study because they are small compared to other astronomical objects. The diffraction limit dictates that resolving the disk of an Earth-sized planet at 10 pc requires a telescope—or array of telescopes—at least 25 km across at optical wavelengths, or 500 km across in the thermal infrared. Short of building a large interferometer to spatially resolve the disks of exoplanets, we must rely on chance and astronomical trickery to map their atmospheres and surfaces.
@@ 2 Mapping Basics
This chapter covers the analysis of planetary light, be it thermal radiation or reflected starlight. Planetary light can be separated from starlight temporally (Charbonneau et al. 2005; Deming et al. 2005), spatially (Marois et al. 2008), or spectrally (Brogi et al. 2012). Transmission mapping is also briefly described below.
@@ 2.1 Transmission Mapping
It is also possible to glean spatial information about an exoplanet’s atmosphere when it transits in front of its host star: transmission mapping, or transit mapping. Transmission mapping shares some concepts with emission and reflection mapping, so is briefly described here. Transmission mapping relies on two geometrical facts: 1) only part of the planetary limb is backlit during transit ingress and egress, and 2) the planetary limb probed in transmission moves on the planet during a transit.
At the start of transit ingress, only the western (leading) limb of the planet is in front of the star and contributes to the transmission spectrum; likewise, the end of egress is only sensitive to the eastern (trailing) limb of the planet. Such transmission mapping has been proposed for detecting the difference in atmospheric scale height (Dobbs-Dixon et al. 2012), winds (Louden and Wheatley 2015), and aerosols (Kempton et al. 2017; Powell et al. 2019) near the eastern and western terminators. Many codes have been developed to perform such transmission mapping (Espinoza and Jones 2021; MacDonald and Lewis 2022; Grant and Wakeford 2023). The ingress-egress asymmetry is most marked for planets that are large compared to their star; the best spatial resolution could be obtained for a giant planet orbiting a white dwarf (e.g., Vanderburg et al. 2020). This phenomenon is conceptually similar to eclipse mapping, but is very sensitive to the limb-darkening of the star. Transmission mapping has already been performed with ground-based high resolution spectroscopy (Ehrenreich et al. 2020). Even in the middle of transit, high resolution data can differentiate between the blue-shifted eastern limb and and the red-shifted western limb (Boucher et al. 2023), analogous to Doppler imaging.
Moreover, the planet will rotate during the transit, i.e., the planetary limb does not exactly coincide with the day–night terminator. In fact, the limb only corresponds to the day–night terminator when the planet is in the center of the stellar disk, which never occurs for a non-zero impact parameter. At the start of transit the limb runs on the dayside of the eastern terminator and on the nightside of the western terminator, and vice versa at the end of transit. The angle between the limb and terminator is proportional to the angular size of the star as seen from the planet, and is therefore most important for ultra short period planets (Nguyen et al. 2020). If a rocky planet has a negligible atmosphere, then topographical features may rotate in and out of view during the transit (McTier and Kipping 2018). Transmission mapping of exoplanets is an exciting growth area, but the remainder of this review focuses on mapping exoplanets using light emitted or reflected by the planets themselves.
@@ 3 Mapping Formalism
Arguably the most important simplification is assuming a static map, ${\partial M}/{\partial t}=0$, as it is generally intractable to map the surface of an unresolved planet when that map is changing. One can map a slowly varying planet, e.g., large-scale cloud patterns on Earth evolve slowly compared to the planet’s rotation (Kawahara and Masuda 2020; Teinturier et al. 2022) and some eccentric planets may still be amenable to eclipse mapping, despite seasonal variations. Beyond these exceptions, one must adopt a parameterized time-variable map, $M(\theta,\phi,t)$. For example, phase curves of eccentric hot Jupiters have been fitted with energy balance models (Lewis et al. 2013; de Wit et al. 2016; Dang et al. 2022). Finally, Cowan et al. (2017) showed that for planets on circular, edge-on orbits one can infer a time-variable map based on the presence of odd harmonics; nonetheless, this is a far cry from mapping the changing surface of a planet.
Barring a changing map, time-variations in flux come in through the kernel. If there are no occultations, then the changing sub-observer and sub-stellar locations dictate the time-variability of the kernel and hence the time-variations in observed flux.
@@ 3.1 The Inverse Problem
If the orientation of a planet is not known a priori, then the problem is even more challenging because the kernel is a function of one or more unknown parameters: spin frequency, spin orientation, etc. Nonetheless, it has been demonstrated in numerical experiments that one can extract a planet’s spin (rotation rate, obliquity and its direction) from reflected lightcurves (Pallé et al. 2008; Oakley and Cash 2009; Kawahara and Fujii 2010, 2011; Fujii and Kawahara 2012; Schwartz et al. 2016; Kawahara 2016; Farr et al. 2018; Nakagawa et al. 2020) or from thermal lightcurves (Gaidos and Williams 2004; Gómez-Leal et al. 2012; Cowan et al. 2012b, 2013). A similar formalism has been developed for imaging exoplanets with a solar gravitational lens (Toth and Turyshev 2023).
Exoplanets can be mapped in monochrome using single-band photometry, or using spectrally-resolved data to produce maps at each wavelength. It is often more insightful, however, to assume that the lightcurves are correlated due to viewing geometry, molecular absorption, or surface spectra. With high spectral resolution, for example, the signal is in the time-variable mean molecular line shape. We point the interested reader to the excellent review of high resolution atmospheric characterization (Birkby 2018) and note that much work remains to be done in extracting spatial information from high resolution data. The inverse problem of mapping a planet using template spectra—let alone retrieving intrinsic surface colors and spectra—based on multi-band photometry or time-resolved spectroscopy is beyond the scope of this review, but has been considered in the literature both in the context of reflected light (Cowan et al. 2009, 2011; Fujii et al. 2010, 2011; Cowan and Strait 2013; Fujii et al. 2017; Kawahara 2020) and thermal emission (Kostov and Apai 2013; Buenzli et al. 2014; Stevenson et al. 2014; Feng et al. 2016; Irwin et al. 2020; Changeat and Al-Refaie 2020; Cubillos et al. 2021).
@@ 4.1 Pixels and Slices
An intuitive approach is to pixelize the planetary surface. The most general pixelization scheme is 2-dimensional, for example, a latitude–longitude grid or Hierarchical Equal Area isoLatitude Pixelization (HEALPix), but more exotic schemes have been developed for occultation mapping (Louden and Kreidberg 2018). Pixels are useful if the kernel moves or changes its shape in more than one direction, e.g., eclipse mapping or rotational reflected light curves at multiple orbital phases.
@@
For rotational mapping, latitudinal information is difficult to extract, so it is often reasonable to consider a map consisting of uniform slices, like a beach ball. This situation includes rotational mapping in the thermal infrared (Knutson et al. 2007) and rotational mapping of reflected light at a fixed orbital phase (Cowan et al. 2009).
@@ 4.3 The Kernel
Due to the unchanging kernel shape for rotational or orbital thermal mapping, a large fraction of spherical harmonics are in the nullspace, limiting the fidelity of retrieved maps. A common strategy is to ignore basis maps in the nullspace, setting their amplitude to zero in fits (Cowan and Agol 2008). But this strategy can sometimes produce retrieved maps that appear to have regions of negative flux, an artifact of the band-limited deconvolution. The solution is to fit for more basis maps—including those in the nullspace: immeasurable fitted parameters may help the map remain everywhere positive (Keating et al. 2019). Such hidden variables can be further constrained by a regularization scheme or Gaussian process. Indeed, the most general and robust approach to mapping is likely a high-order orthogonal map expansion—many spherical harmonics or pixels—embedded in a hierarchical model (Farr et al. 2018; Luger et al. 2021). If the prospect of fitting unseen and unknowable parameters by relying on priors is daunting, rest assured that you can often obtain a good fit and reasonable map by adopting an altogether different parameterization, such as slices or pixels (Beatty et al. 2019).
@@
If a planet passes directly behind its star, then one may map its day side. Eclipse mapping can be applied to either planetary emission or reflected light, but the contrast ratio makes the latter daunting. While the eclipse depth is proportional to the integrated day side brightness of the planet, the detailed shape of ingress and egress is a function of the spatial distribution of flux on the planet’s day side (Figure 3). Although eclipses of planets by their host star have so far garnered the most attention, occultations by moons or other planets are also possible, especially in packed planetary systems like TRAPPIST-1 (Veras and Breedt 2017; Luger et al. 2017).
@@ 5 Current Results & Future Prospects
The atmospheres of a few dozen exoplanets have been mapped via their thermal phase curves, e.g., the ExoAtmospheres and exoplanet.edu databases. These phase measurements has so far been mostly IRAC photometry from the Spitzer Space Telescope (e.g., Bell et al. 2021; May et al. 2022), or spectroscopy with the Hubble Space Telescope’s WFC3 (e.g., Stevenson et al. 2014; Kreidberg et al. 2018). We send the interested reader to the brief history of thermal phase curves in Charnay et al. (2022). Phase measurements from the James Webb Space Telescope have started being published, but so far focusing primarily on photometry (Mikal-Evans et al. 2023; Bell et al. 2023; Kempton et al. 2023). Rotational phase variations of brown dwarfs have also been measured and converted into brightness maps (e.g., Apai et al. 2017, and chapter by Artigau).
The Ariel mission could measure spectroscopic phase curves for dozens—if not hundreds—of planets and brown dwarfs (Charnay et al. 2022). The wavelength coverage of Ariel and JWST should be able to discriminate between thermal emission and reflected light (Schwartz and Cowan 2015; Parmentier et al. 2016; Keating and Cowan 2017), as well as disentangling phase variations, star spots, Doppler beaming, ellipsoidal variations, and gas accretion (Cowan et al. 2012a; Knutson et al. 2012; Esteves et al. 2015; Dang et al. 2018; Bell et al. 2019).
@@
Ground-based direct imaging could detect variations in thermal emission from directly-imaged planets due to rotation of patchy clouds in and out of view (Biller et al. 2021), as is currently performed for brown dwarfs. The next generation of ground based telescopes should enable both photometric mapping and Doppler tomography of giant planets (Crossfield et al. 2014), while the Large Interferometer for Exoplanets (Quanz et al. 2022) could map the climates of temperate terrestrial planets (e.g., Gaidos and Williams 2004; Cowan et al. 2012c; Mettler et al. 2023).
@@ Acknowledgements.
  • “Detecting Habitability”
  • “Characterization of Exoplanets: Secondary Eclipses”
  • “Characterization of Exoplanets: Observations and Modeling of Orbital Phase Curves”
  • “Variability of Brown Dwarfs”
  • “Spectroscopic Direct Detection of Exoplanets”